# SeeThruFinger: See and Grasp Anything with a Multi-Modal Soft Touch

### v3

@@ Abstract
We present SeeThruFinger, a Vision-Based Tactile Sensing (VBTS) architecture using a markerless See-Thru-Network. It achieves simultaneous visual perception and tactile sensing while providing omni-directional, adaptive grasping for manipulation. Multi-modal perception of intrinsic and extrinsic interactions is critical in building intelligent robots that learn. Instead of adding various sensors for different modalities, a preferred solution is to integrate them into one elegant and coherent design, which is a challenging task. This study leverages the in-finger vision to inpaint occluded regions of the external environment, achieving coherent scene reconstruction for visual perception. By tracking real-time segmentation of the Soft Polyhedral Network’s large-scale deformation, we achieved real-time markerless tactile sensing of 6D forces and torques. We demonstrate the capable performances of the SeeThruFinger for reactive grasping without using external cameras or dedicated force and torque sensors on the fingertips. Using the inpainted scene and the deformation mask, we further demonstrate the multi-modal performance of the SeeThruFinger architecture to simultaneously achieve various capabilities, including but not limited to scene inpainting, object detection, depth sensing, scene segmentation, masked deformation tracking, 6D force-and-torque sensing, and contact event detection, all within a single input from the in-finger vision of the See-Thru-Network in a markerless way. All codes are available at https://github.com/ancorasir/SeeThruFinger.
Keywords In-finger Vision Multi-Modal Perception Segmentation Tracking Scene Inpainting
@@ 1 Introduction
Vision-based sensing is probably the most widely adopted perceptual modality in modern robotics, transforming the unstructured environment into a structured embodiment for reasoning and learning in an object-centric manner [1]. Besides bioinspiration from the human’s hand-eye system, robotic vision provides a unified representation of the 3D world into matrices of 2D pixels that have evolved rapidly in both theoretical foundations and engineering applications, which is further accelerated by recent advancements in data science and deep learning [2]. While the classical methods focus on the visual features to extract reusable information analytically [3], recent development exploits data-driven methods by systematically integrating promptable segmentation tasks, pre-trained foundation models, and web-scale datasets to segment anything [4]. Such capability provides researchers with further insights to achieve object tracking [5], object classification [6], pose estimation [7], occlusion inpainting [8], and scene reconstruction [9] with pixel-level details in an automatic workflow, paving the path towards a multi-modal physical embodiment of machine intelligence for robot learning [10]. It should be noted that vision-based perception remains a leading choice for many robot learning research and applications despite the growing accessibility and availability of other perceptual modalities.
Due to basic principles such as line-of-sight, ambient lighting, and surface reflection, problems such as occlusion remain challenging in robotic vision. Occlusion is a practical problem caused by optical physics, which is inevitable for robot workstations that mainly adopt cameras as the only or primary sensor for extrinsic perception. To perform iterative visual verification during object manipulation [11], researchers usually need to add more cameras [12] or move the gripper to a particular pose to expose the fingers in the viewing range for grasp result checking. These solutions make trade-offs in hardware, integration, or time cost. While it is always possible to add more sensors or use more powerful ones, it remains a difficult design task to integrate multiple sensory modalities within a compact form factor for robotic manipulation [13].
1.1 Occlusion in Common Hand-Eye Systems
The robotic manipulation problem is typically set up as a Hand-Eye system [14], as illustrated in Fig. 1A. The “Hand” represents a gripper attached to an “Arm” manipulator’s flange for handling physical contact with objects. The “Eye” is an external camera for visual perception. Since only light signals with unobstructed lines of sight could be captured by the camera, occlusions commonly occur when some objects are stacked on the other. For example, Object 1 in Fig. 1A is placed at the top with its complete shape captured in the image, but parts of Objects 2 and 3 are occluded since they are placed underneath. Inspired by the human approach, roboticists have investigated different solutions to address occlusion via better tools, more skills, or multi-modal interactions.
Human usually adopts the better-tools solution by upgrading hardware or software without disrupting the original task pipeline. For example, in many scenarios, the occlusion problem could be partially alleviated by using more capable machine vision systems, such as 3D vision capturing the scene in dense point cloud data [15], or more efficient learning algorithms to handle complexities in feature extraction [16].
The more-skills solution is also a human-inspired strategy, which involves rearranging the scene via added skills in motion planning so that the occluded objects can be visually captured in sufficient detail. Motion planning using convex optimization enables added skills for robotic arms to move around obstacles for object manipulation [17]. Sampling-based planning with hierarchical graphs also provides capabilities to retrieve objects in cluttered scenes by moving obstacles around until the target object is found [18].
The multi-modal solution introduces further perceptual modalities to aid in handling uncertainties caused by occlusions. Tactile sensory from the skin, especially those on the fingertips, is a biological answer to this problem with evolutionary evidence [19]. Introducing tactile sensing on top of robotic vision enhances learning manipulation skills, formulating a coarse-to-fine hierarchy between globalized scene understanding and localized manipulation dynamics [20].
1.2 Vision-based Tactile Sensing Architectures
The robotics research community is interested in Vision-Based Tactile Sensing (VBTS) to complement visual perception through contact [21, 22]. While researchers have developed various tactile sensors utilizing different sensing principles [23, 24], VBTS solutions remain preferred among robotic researchers in academia and applications [25]. Some advantages include easy access to camera sensors, unified data structure, shared theories and learning models from computer vision, and deep learning. However, due to the size of the camera board, the field of view, choice of lenses, and electrical wiring, it remains a design problem for vision-based tactile sensors to be seamlessly compatible with various multi-fingered gripper systems. Recent literature reports several in-depth review articles on the technical specifications and theoretical foundations of VBTS [26]. In the following, we focus our review on the design analysis of two VBTS architectures, one with an Illuminated Chamber in Fig. 1B and another with a See-Thru-Network in Fig. 1C.
In-finger vision represents a general architecture of VBTS, which mainly features an Illuminated Chamber design illustrated in Fig. 1B. A miniature camera is installed within a closed chamber, physically isolated from the ambient environment. Therefore, active lighting using LEDs is usually required to enable the proper functions of the camera [27]. In front of the lens is a soft membrane made from transparent elastomer, covered with a layer of opaque coating on the outer surface, blocking the line of sight. Before contact occurs on the soft membrane’s outer surface, VBTS with an Illuminated Chamber design usually cannot capture visual features of the external environment. Instead, it requires a touch event on the soft membrane to induce imprinted textures, assisted by active lighting in a predesigned color sequence inside to enable visual capturing of the refined textures via photometric processing. In addition, researchers usually paint artificial marker arrays on the inner surface of the soft membrane [28] or even directly embed these markers inside [29] to enable estimation of the tactile forces and torques. A representative application of VBTS with an Illuminated Chamber design is as a profiler for surface measurement by moving the sensor directly toward the target surface. However, like human fingers, the robotic finger usually detects tactile interactions on the side of the fingertip when the gripper closes, assisted by the manipulator’s general motion to approach the object, as shown in Fig. 2A.
As a result, researchers usually need to rotate the sensor during implementation, as shown in Fig. 2B, when fixing VBTS with an Illuminated Chamber on the robotic fingertip. Researchers recently explored design modifications by adding a reflective and opaque mirror inside to change the light path, as shown in Fig. 2C(i), so the sensor could be more conveniently fitted on robotic fingertips [30]. Design modification could be done using a translucent coating instead so that partial light from the outside could pass through the soft membrane [31, 32]. This enables visual perception of the scene at a close distance before the contact event, as the image is usually blurry due to the translucent coating used. A similar design modification is to use a reflective in Fig. 2C(i) or a translucent mirror in Fig. 2C(ii) to achieve similar purposes.
Recent research also reports another architecture for VBTS using a See-Thru-Network, as illustrated in Fig. 1C. A Soft Polyhedral Network (SPN) replaces the soft membrane with a hollow space inside to enable a see-through design [33, 34]. The SPN is a soft robotic metamaterial capable of geometric adaptations in omni-directional interactions. One can fix fiducial markers inside the SPN, transforming its spatial adaptations into the marker’s 6D movement, visually tracked by the miniature camera inside. This enables an open base frame, allowing the camera lens to be fully exposed to the external environment [35]. Leveraging the See-Thru-Network design, the images captured contain two layers of information following the line-of-sight principle. The first layer is the structural state of the SPN, followed by a second layer showing a partial view of the scene and objects occluded by the SPN. When adopting a finger-like design for the SPN, primary interactions occur on the primary side of the SPN, allowing the whole finger to be installed with the in-finger camera directly facing the scene. As the manipulator approaches the object, an occluded view of the scene and objects is always visible to the in-finger camera but impeded by a fixed image of the SPN. When the gripper closes in a lateral direction until contact begins, the SPN interacts with the object through omni-directional adaptation, fully captured by the in-finger vision.
While both VBTS architectures feature an in-finger vision component, they provide tactile perception at different levels of detail when integrated with a robot manipulation system. Using photometric imaging within an enclosed chamber, the Illuminated Chamber design excels in refined texture reconstruction within a small range of soft membrane deformation. On the other hand, See-Thru-Network provides large-scale geometric adaptation while producing high-performing 6D force and torque estimation that is not limited to common on-land scenarios but applies to the underwater environment [36, 37].
1.3 Proposed Method and List of Contributions
This study proposes the SeeThruFinger architecture, as shown in Fig. 3, using the Soft Polyhedral Network for simultaneous visual perception of the scene and tactile sensing of physical contacts via in-finger vision, achieving markerless Vision-Based Tactile Sensing with a See-Thru-Network. Before the contact begins, a static mask of the SPN is used to inpaint the occluded scene from in-finger vision via a head-up movement. The inpainted in-finger vision reconstructs the scene, enabling typical downstream capabilities such as object detection for grasping. Then, we implement markerless, real-time, and large-scale deformation tracking of the SPN via video object segmentation using the SPN’s static mask. Real-time monitoring of the SPN’s deformation mask enables us to achieve markerless tactile sensing by learning to infer high-performing 6D forces and torques during grasping. Finally, we demonstrate the feasibility of our proposed SeeThruFinger architecture through a reactive grasping task. We also demonstrate SeeThruFinger’s potential in multi-modal perceptual embodiment to simultaneously achieve various downstream capabilities such as object detection, depth sensing, and scene segmentation after scene inpainting, which benefits from the growing adoption of foundational models. Overall, the SeeThruFinger architecture holds the potential to systematically simplify the hardware requirement of a typical Hand-Eye system through multi-modal perceptual fusion using a single in-finger vision sensor. Contributions of this study are listed as follows:
This study proposes the SeeThruFinger architecture for Vision-based Tactile Sensing (VBTS), achieving simultaneous decoupling of tactile sensing for physical contact and multi-modal visual perception, such as object detection, depth sensing, and scene segmentation, within a single in-finger vision system.
This study is among the first to achieve markerless VBTS using real-time masked segmentation and scene inpainting by leveraging the See-Thru-Network’s occluded design features for enhanced multi-modal tactile perception.
This study also validates the proposed method’s competitive performance against traditional hand-eye systems in a multi-object reactive grasping demonstration via a single SeeThruFinger without using external cameras, dedicated tactile sensors, or force-torque sensors.
2 Preliminaries & Problem Formulation
@@
We begin by denoting an image captured by a camera at timestep . The and are the user-defined output settings for the image size, measured in the total count of pixel rows and columns. Depending on the camera selection, the th row and th column pixel entry of the image may contain the brightness intensity of a single channel for grayscale images () or three channels in red, green, and blue (RGB) for color images (). Each pixel entry usually stores information in an 8-bit format, offering 256 conditions ranging from 0 to 255. Furthermore, we can add another data channel () to the RGB image to include each pixel’s transparency state.
2.1 Markerless Design of the SeeThruFinger
@@
Shown in Fig. 4 is the Hand-Eye system adopted in this study, built with aluminum extrusion based on the DeepClaw workstation [14], housing a collaborative robot (UR10e from Universal Robots) on a pedestal and a tabletop in front covered by black fabrics. Although an “Eye” camera is fixed on the post shown in Fig. 4A, it is intentionally unplugged (therefore with a strikethrough) as we intend to implement visual perception of the scene using in-finger vision only from the SeeThruFinger. Figure 4B shows an enlarged view of the “Hand,” a two-fingered gripper (Model AG-160-95 from DH-Robotics) with its fingertips replaced by the SeeThruFingers. Each SeeThruFinger contains a soft, omni-adaptive, markerless finger based on a variant of the Soft Polyhedral Network designs [34], mounted on a camera housing 3D-printed by UV-curable transparent resin (Somos WaterShed XC 1112), where a miniature camera (Chengyue WX605 from Weixinshijie) is fixed inside, as shown in Fig. 4C.
@@
One can customize the base plate to install the SeeThruFinger on Force-Torque (FT) sensors for testing or on a gripper as fingertips for soft and adaptive grasping. Figure 4D shows the objects used for testing in this study, including a Pringles Can from the Yale-Columbia-Berkley (YCB) object sets [38] and others commonly available at research labs, including spray can, mineral water bottles filled halfway with water, medicine bottle, Bosche electrical screwdriver, pouch cell. Although both SeeThruFingers on the gripper are functional, we only used the one whose viewing range partially covers the table at every frame throughout the following experiments.
2.2 Formalizing Multi-Modal VBTS via a See-Thru-Network

### v4

# Multi-Layered Reasoning from a Single Viewpoint for Learning See-Through Grasping

@@ Abstract
Sensory substitution enables biological systems to perceive stimuli typically obtained by another organ, which is inspirational for physical agents. Multi-modal perception of intrinsic and extrinsic interactions is critical in building an intelligent robot that learns. This study presents a Vision-based See-Through Perception (VBSeeThruP) architecture that simultaneously perceives multiple intrinsic and extrinsic modalities via a single visual input in a markerless way, all packed within a soft robotic finger using the Soft Polyhedral Network design. It is generally applicable to miniature vision systems placed underneath deformable networks with a see-through design, capturing real-time images of the network’s physical interactions induced by contact-based events overlayed on top of the visual scene of the external environment, as demonstrated in the ablation study. We present the VBSeeThruP’s capability for learning reactive grasping without using external cameras or dedicated force and torque sensors on the fingertips. Using the inpainted scene and the deformation mask, we further demonstrate the multi-modal performance of the VBSeeThruP architecture to simultaneously achieve various perceptions, including but not limited to scene inpainting, object detection, depth sensing, scene segmentation, masked deformation tracking, 6D force/torque sensing, and contact event detection, all within a single sensory input from the in-finger vision markerlessly.
Keywords Multi-Modal Learning Latent Representation Tactile Robotics
@@ 1 Introduction
Sensory substitution [1] is a technique where one sensory modality is used to provide information typically perceived by another. Biological systems utilize visual cues to infer contact-based information, serving as inspiration for achieving multi-modal robotic perception [2]. In classical settings, such as the “Rubber Hand Experiment” [3], research has shown that humans can translate contact-based sensations into visual or auditory signals, enabling them to see through touch [4] or hear textures [5]. This concept naturally extends to understanding deformable interactions in both human and biological systems, where visual observations of deformations can infer underlying contact-based interactions and material properties [6]. This principle forms the basis of Vision-Based Deformable Perception (VBDeformP), particularly in robotic manipulation, where the growing adoption of miniature vision sensors [7, 8] and robot learning methods [9, 10] are also notable. Accurately perceiving contact-based information is crucial for physical agents, such as robots, to effectively grasp and manipulate objects [11]. Leveraging visual cues of deformations offers a promising solution [12].
Robotics research has increasingly explored vision-based techniques to understand deformable objects [13] and their interactions with the object-centric environment [14]. Simultaneously achieving a comprehensive perception of the scene (i.e., object detection, recognition, and spatial understanding) and the touch-based interactions (i.e., surface texture, material properties, contact forces) during object manipulation with robots is an active research area [10].
However, achieving multi-modal perception within a single, compact visual input, especially when the sensor is embedded inside a soft robotic fingertip, poses a significant challenge [15]. The deformable interface between a miniature camera and the external environment plays a critical role in the accuracy and applicability of the system. Its design directly impacts the quality of both visual and contact-based information acquired, making it an emerging research area in tactile robotics [16]. An ideal system would be able to decouple the contact-based interaction gathered from the soft interface’s deformation from the miniature vision that also perceives the scene, all from a single visual input.
This work presents a Vision-based See-Through Perception (VBSeeThruP) architecture that enables multi-layered reasoning to learn reactive grasping from a single visual input behind a deformable network, serving as a markerless representation of fingertips, as shown in Fig. 1. The VBSeeThruP approach addresses the challenges of concurrently perceiving both finger deformation induced by contact-based interactions and rigid objects within the scene, all within a unified visual input for physical agents. For enhanced interactions with omni-directional adaptation of objects, we adopted the Soft Polyhedral Network (SPN) reported in recent research [17, 18], and further achieved markerless representations through multi-modal perception. The miniature vision captures real-time images of the SPN’s physical deformation during contact-based interaction while preserving partial information of the external scene. Before contact, a static SPN mask is used to inpaint the occluded scene via a brief “head-up” camera movement. Subsequently, video object segmentation of the SPN’s large-scale deformations provides markerless contact-based perception by estimating 6D forces and torques during contact. We demonstrate the viability of this approach through an ablation study and a reactive grasping task, further showcasing its potential to perform object detection, depth sensing, and scene segmentation once occluded regions are inpainted. This offers a unified solution that could streamline traditional hand-eye systems into a single in-finger vision sensor.
2 Related Works
Vision underpins robotic perception by enabling object detection, localization, and scene understanding [19, 20]. However, properties such as stiffness, friction, and mass are difficult to infer from vision alone [21, 22], calling for integration of alternative modalities.
2.1 Differentiating Vision-based Sensing and Perception
In robotics and cognitive science, sensing and perception are closely related but fundamentally distinct concepts, as summarized in Table 1.
| Aspect | Sensing | Perception |
| Data Type | Raw, unprocessed signals | Processed, interpreted information |
| Role | Data collection | Interpretation, understanding, and decision-making |
| Implementation | Primarily hardware-level | Primarily computational or cognitive algorithms |
| Complexity | Low-level, simpler | High-level, complex, involving learning or inference |
| Outcome | Raw sensor measurements | Meaningful representations or models |
| Example (Visual) | Camera pixels | Object detection, scene understanding |
| Example (Tactile) | Force/pressure readings | Recognizing object stiffness or shape |
Sensing provides the foundational information required by robots or intelligent agents to interact with the physical world. It refers to directly acquiring raw data from the environment using sensors. It is the preliminary stage of gathering signals or measurements. It typically involves hardware (e.g., cameras, microphones, force sensors, temperature sensors); produces unprocessed, quantitative, often noisy data; and measures passive or active physical signals without interpreting their meaning. For example, a camera captures pixel-level images, or a microphone records sound waves, which are typically considered forms of sensing. Perception, on the other hand, builds on sensing data, enabling robots to make informed decisions, adapt to new situations, and interact intelligently with the environment or humans. It involves processing, interpreting, and understanding sensory data to derive meaningful information or actionable insights about the environment or state. It often refers to the computational or cognitive processes involved in analyzing and interpreting sensor data, which include filtering, classification, recognition, inference, and decision-making tasks. Perception bridges the gap between raw sensing and actionable knowledge or behavior. For example, identifying objects within an image, recognizing human speech from audio signals, or estimating forces from tactile sensor data are generally considered forms of perception.
2.2 A General Classification of Vision-based Perception
An object-centric representation [23] leads to rigid-body assumptions or deformable interactions for visual understanding and reasoning, promoting robotic vision via Vision-based Rigid Perception (VBRigidP) and Vision-based Deformable Perception (VBDeformP), which are the leading approaches.
2.2.1 Vision-based Rigid Perception (VBRigidP)
VBRigidP methods traditionally assume rigid objects, leveraging tools such as Faster R-CNN [24] or YOLO [25] for detection, combined with pose estimation techniques (PnP, ICP) [26, 27] for manipulation tasks [28]. Although effective in structured industrial settings, VBRigidP and the rigid-object assumption become limiting for real-world scenarios involving deformable or partially deformable objects.
2.2.2 Vision-based Deformable Perception (VBDeformP)
VBDeformP extends beyond rigid bodies by tracking non-rigid shapes [29, 30]. Soft robotics, cloth manipulation, and surgical robotics applications exploit visual deformation cues to infer forces or material properties [31, 32]. By analyzing deformation in video, robots can estimate contact forces [33], plan tasks for large-scale cloth handling [31], or reconstruct object geometry [34], all without the need for specialized tactile sensors.
Emerging from VBDeformP is Vision-based Tactile Perception, which interprets tactile interaction by analyzing deformation images [35, 21]. Research has shown that detailed force and contact geometry can be inferred from images of a soft interface undergoing compression [36, 35], which can be further embedded into multi-fingered systems for object manipulation [37, 38]. Recent research has demonstrated tactile perception inferred at the sensor (e.g., normal and tangential forces [35], vibration [39], thermal [21], pretouch proximity [40]), contact (e.g., contact geometry [35], force and torque [18], contact events [41], material properties [42]), object (e.g., object localization [43], shape [33, 44], mass and dynamics [36], contents of containers [45]), and action (e.g., can be inferred from previous perceptual modalities for action selection and initialization, tactile feedback for low-level control, action termination, action outcome detection, and action outcome verification) levels [46], providing a cost-effective solution for tactile robotics [47].
It should be noted that the Vision-based Tactile Perception mentioned above is typically referred to as Vision-based Tactile Sensing (VBTS) in the computer vision and robotics literature. However, such terminology has caused some confusion among researchers in the cognitive science field, as the raw output signals obtained are camera pixels, and the subsequent tactile information is computationally processed to infer contact-based meanings for object manipulation. In addition, it could be argued that the tactile data collected using the VBTS method is not directly measured from compression readings regarding force or pressure on the interaction surface, but rather processed algorithmically based on the raw image pixels, making it more of a perception method than a sensing method. Without losing generality, we describe this approach as a subcategory of the VBDeformP method in this work, following the general review elaborated in Sec. 2.1.
2.2.3 Vision-based See-Through Perception (VBSeeThruP)
In this work, we introduce Vision-based See-Through Perception (VBSeeThruP), a novel approach that combines the benefits of both VBRigidP and VBDeformP to achieve multi-modal perception from a single visual input by reasoning about information obtained from different image layers during in-finger imaging. Cognitive science research on sensory substitution inspired our research in VBSeeThruP, as we can use a single visual sensing to infer multiple perceptual modalities by reasoning between various layers of information the camera captures. For example, our proposed VBSeeThruP architecture involves inferring contact-related information by tracking masked deformation from the See-Thru-Network (which is not limited to the Soft Polyhedral Network design exemplified in this work). The VBSeeThruP can also infer visual-related information after scene inpainting and the use of foundational models for a rich set of perceptual modalities in an object-centric scene. The combined effect is not achievable using VBRigidP or VBDeformP methods alone, and the resultant effect is enabled by introducing a See-Thru-Network in front of the in-finger camera. For example, when the network structure is 100% filled without see-through capabilities, it becomes equivalent to a VBTactileP (or VBTS) system. When it is 0% filled with complete see-through capabilities, it becomes equivalent to a VBRigidP system.
2.3 Multi-Modal Perception in Robotics with Vision
Combining vision and touch provides richer representations, improving classification [48], pose estimation [49], and in-hand manipulation [50]. This fusion is instrumental in cluttered or occluded environments [51, 52]. Pairing visual and auditory cues boosts performance in human action classification and material recognition [53, 54]. Multi-sensory attention mechanisms (vision, audio, touch) enhance performance for complex activities, such as pouring or packing [55, 56]. Recent multi-modal large language models (LLMs) combine visual perception with language-based reasoning, enabling zero-shot and few-shot policy learning [57, 58]. These systems exhibit stronger physical reasoning when additional modalities are incorporated [59, 60], widening opportunities for flexible, data-driven manipulation [61].
3 Proposed Methods
@@
Our approach begins by denoting an image captured by a camera at timestep . The and are the user-defined output settings for the image size, measured in the total count of pixel rows and columns. Depending on the camera selection, the th row and th column pixel entry of the image may contain the brightness intensity of a single channel for grayscale images () or three channels in red, green, and blue (RGB) for color images (). Each pixel entry usually stores information in an 8-bit format, offering 256 conditions ranging from 0 to 255. Furthermore, we can add another data channel () to the RGB image to include the transparency state of each pixel.
3.1 Markerless Design of the See-Thru-Network
Figure 2 is the Hand-Eye system adopted in this study, built with aluminum extrusion based on the DeepClaw workstation [62], housing a collaborative robot (UR10e from Universal Robots) on a pedestal and a tabletop in front covered by black fabrics.
@@ 3.1 Markerless Design of the See-Thru-Network
Although an “Eye” camera is fixed on the post shown in Fig. 2A, it is intentionally unplugged (therefore with a strikethrough) as we intend to implement visual perception of the scene using in-finger vision only behind a soft fingertip featuring a See-Through design using deformable materials. Figure 2B shows an enlarged view of the “Hand,” a two-fingered gripper (Model AG-160-95 from DH-Robotics) with its rigid fingertips replaced by the soft ones. Each soft fingertip contains a soft, omni-adaptive, markerless structure based on a variant of the Soft Polyhedral Network designs [63], mounted on a camera housing 3D-printed by UV-curable transparent resin (Somos WaterShed XC1112), where a miniature camera (Chengyue WX605 from Weixinshijie) is fixed inside.
@@
As shown in Fig. 2C, the base plate can be customized to accommodate the installation of soft fingertips on force/torque (FT) sensors for testing or to use a gripper as a fingertip for soft and adaptive grasping. Figure 2D shows the objects used for testing in this study, including a Pringles Can from the Yale-Columbia-Berkley (YCB) object sets [64] and others commonly available at research labs, including a spray can, mineral water bottles filled halfway with water, a medicine bottle, a Bosche electrical screwdriver, pouch cell. Although both soft fingertips are functional, we only used the one whose viewing range partially covers the table at every frame throughout the following experiments.
3.2 Formalizing Multi-Modal VBSeeThruP
For an image taken by an in-finger vision camera, the resultant image is essentially the same as the general representation reproduced in Fig. 3, with the ambient environment in gray.